How long would it take a $3000 investment to double if it is invested at 7% compounded semiannually. Rounded to the nearest tenth of a year.
Answer by stanbon(75887)

How long would it take a $3000 investment to double if it is invested at 7% compounded semiannually. Rounded to the nearest tenth of a year.
---------------------------
A(t) = P(1+(r/n))^(nt)
6000 = 3000(1+(0.07/2))^(2t)
2 = (1.035)^2t
(2t)log(1.035) = log(2)
2t = 20.1488
t = 10.07 years
======================
Cheers,
Stan H.